OpenStudy (anonymous):

The sum of two consecutive integers is at least 36. What is the least possible pair of integers? A. 16 and 17 B. 17 and 18 C. 18 and 19 D. 19 and 20

OpenStudy (whpalmer4):

How many of those pairs add up to at least 36? If more than one does, which one is the smallest?

OpenStudy (anonymous):

i say B

OpenStudy (anonymous):

Try them all out if you have doubts.

OpenStudy (whpalmer4):

Does 17 + 18 = at least 36?

OpenStudy (anonymous):

@Luis_Rivera, they have to be integers.

OpenStudy (whpalmer4):

@Luis_Rivera check your arithmetic

OpenStudy (anonymous):

im confused

OpenStudy (whpalmer4):

@pinkpearl12000 you have 4 pairs of numbers. you're looking for the smallest pair that adds up to at least 36. Add up each pair, pick the smallest one that adds up to at least 36. what is confusing?

OpenStudy (whpalmer4):

@Luis_Rivera Dude, you aren't reading the problem statement correctly.

OpenStudy (anonymous):

So try all your pairs: 16+17 = 33 17+18 = 35 18+19 = 37 19+20 = 39 So The first two ones don't add up to 36, so they are out. Next you are left with 18+19=37 and 19+20=39 So 37<39 That means 18+19 < 19+20 So the smallest pair of integers that add up to at least 36 is 18+19.

OpenStudy (anonymous):

Ohhhh,but that adds up to 37

OpenStudy (whpalmer4):

37 is at least 36, isn't it?

OpenStudy (whpalmer4):

No, you didn't execute any of your methods correctly.

OpenStudy (whpalmer4):

And only the first one would give you the right answer.

OpenStudy (anonymous):

thanks you lol

OpenStudy (anonymous):

C. 18 and 19